OpenStudy (anonymous):

Given the function f(x) = 6x2 −13, what is f(-3)?

OpenStudy (anonymous):

what do you get when you plug in -3 for X in that function? 6(-3)^2 - 13 =

OpenStudy (sburchette):

For this, substitute -3 for every x in the function.\[f(-3)=6(-3)^2-13\]\[f(-3)=6(9)-13\]\[f(-3)=54-13\]\[f(-3)=41\]

OpenStudy (anonymous):

thank you!!!!!!!